If 730-nm and 640-nm light passes through two slits 0.63 mm apart, how far apart are the second-order fringes for these two wavelengths on a screen 1.2 m away?
Answer #1

We have relation for distance of second bright fringe from center fring

yn=(D\timesn)\times(2\lambda/d)

Here

y2=2D\lambda1/d (Taking n=2)

y2,1= 2D\lambda2/d

so separation is

y2-y2,1=(2D\lambda1/d) - (2D\lambda2/d)

= 2D (\lambda1-\lambda2)/d

putting all vaules

= 2\times1.2 (730\times10-9-640\times10-9)/(0.63\times10-3)

= 3.43\times10-4 m ANS
Answer #1

We have relation for distance of second bright fringe from center fring

yn=(D\timesn)\times(2\lambda/d)

Here

y2=2D\lambda1/d (Taking n=2)

y2,1= 2D\lambda2/d

so separation is

y2-y2,1=(2D\lambda1/d) - (2D\lambda2/d)

= 2D (\lambda1-\lambda2)/d

putting all vaules

= 2\times1.2 (730\times10-9-640\times10-9)/(0.63\times10-3)

= 3.43\times10-4 m ANS